OpenStudy (anonymous):

determine if (3/4,-2) is a solution to -4x-7y please show all work thanks

OpenStudy (anonymous):

are you missing anything in the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3/4,-2 -4x-7y=11

OpenStudy (anonymous):

-4(3/4)-7(-2)=11 -3+14=11 11=11 is a solution